OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y

OpenStudy (anonymous):

to get the total votes y, multiply every term by 4

OpenStudy (anonymous):

can u message me and tell me the answer please cuz I don't get it

OpenStudy (anonymous):

do you have problem to understand how they got the formula? x - 62,500 = 0.25y

OpenStudy (anonymous):

no I need the answer please

OpenStudy (anonymous):

x - 62,500 = 0.25y Maya - 62,500 = sally ok so maya has 62,500 more votes than Sally sally has 25% of the votes Maya has 75% of the votes Maya has 62,500 votes more Maya has 50% more 62,500 ~50% Maya=75% -62,500=50% = 25% total votes y=62,500 * 2 = 125000 =100% Maya had 0.75 x 125000 =93750 votes Sally had 0.25 x 125000 = 31250 votes let's check: Maya - 62,500 = Sally 93750 - 62,500 = Sally 93750 - 62,500 = 31250 31250 = 31250.

OpenStudy (anonymous):

x - 62,500 = 0.25y x = 0.75 y <-- this one was added because they are finalists! in a final 25% sally plus Maya must be 100% so Maya is 75% ! 0.75 y - 62,500 = 0.25y ---subtract 0.25 y from both sides 0.5 y -62,500 = 0 ---add 62,500 .5 y = 62,500 ---multiply by 2 y = 125,000